OpenStudy (anonymous):

How would you solve: Z^2+2z+1 * 5/z+1?

OpenStudy (anonymous):

Is this about complex numbers?

OpenStudy (anonymous):

Lol no...it's multiplying rational expressions.

OpenStudy (anonymous):

If that is the case, then multiply the fraction by a fraction that has both numerator and denominator with the opposite sign of the expression of the denominator of the first fraction. That is:\[\frac{ z^2+2z+1*5 }{z+1}*\frac{ z-1 }{ z-1}=\frac{ (z^2+2z+5)(z-1) }{ 1^2+1^2 }\] You can carry on from here.

OpenStudy (anonymous):

Ok. Thank you so much!

OpenStudy (anonymous):

No problem.